Circular No. 8663

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

HD 109962 F. Walter, Stony Brook University; H. E. Bond, Space Telescope Sci-

ence Institute; and A. Pasten, Cerro Tololo Interamerican Observatory, report that a spectrum of HD 109962 (= NSV 19448), obtained with the SMARTS 1.5-m telescope on Jan. 16.2 UT, shows He II 468.6-nm emission (equivalent width −0.29 nm) and possible weak emission at the 464.0-nm C III/N III feature (equivalent width −0.12 nm) superposed on the star’s F2 V spectrum. In a second spectrogram obtained on Jan. 19.3, however, the high-excitation emission is absent. As described by S.

Otero (http://ar.geocities.com/varsao/NSV 19448.htm), HD 109962

is an eclipsing binary with a period of 0.89 days that additionally shows outbursts of ∼ 0.8 mag and duration ∼ 40 days occurring about once a year, according to ASAS-3 data, with the latest outburst having begun in late December 2005. The variable He II emission suggests a close compact companion undergoing a transient accretion episode, or one that was eclipsed during the second spectrogram. Observations at all wavelengths, including x-ray and ultraviolet, are encouraged.

COMET C/2006 A2 (CATALINA) E. Christensen, Lunar and Planetary Laboratory, writes that the dis-

covery and follow-up images from the Catalina Schmidt telescope do not look cometary, though the object’s images obtained by E. C. Beshore from the Mt. Lemmon 1.5-m reflector on Jan. 24 display a round 8′′ coma in four 60-s stacked images, with no visible tail.

This comet’s designation was inadvertently assigned to the wrong halfmonth (IAUC 8662), but the designation will remain as given.

COMET 73P/SCHWASSMANN-WACHMANN CCD observations of the secondary component to comet 73P (cf. IAUC

8659, 8660) obtained by R. Hill with the Catalina 0.68-m Schmidt telescope (60-s exposures) on Jan. 23.5 UT show a tail ∼ 40′′ long in p.a. 290^{o}. The available astrometry of this component, together with orbital elements (computed assuming identity with component B observed in 1995 and 1996, after Sekanina 2005, ICQ 27, 229) and an ephemeris, were published on MPEC 2006-B27.

2006 January 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green